I am allowed to chose a method of choice...I am primarily confused by the way the equation is written. y + 2 = 0
5x - 2y= 9
Not sure how to use the substitution method when the "y" does not start with y=
I am also not sure how to use addition method when there is not an "x" in the first equation.
Thank you very much for your time and help in advance!!

Answer by mananth(16946)

y+2=0
-2 -2
Y=-2
SUBSTITUTE IN SECOND EQUATION
5x-2y=9
5x-2(-2)=9
5x+4=9
-4 -4
5x=5
/5
x=1
------------------
0x+1y=2...................(1)
5x-2y=9....................(2)
multiply eq.(1) by -2
we get
0+2y=-4
5x-2y=9
add up
5x=5
/5
x=1
y=2